# When Should Selfish Miners Double-Spend?

### v1

@@ Abstract
Although, both double-spending and selfish-mining attacks have been extensively studied since the “Bitcoin” whitepaper of Nakamoto and the “majority is not enough” paper of Eyal and Sirer, there has been no rigorous stochastic analysis of an attack that combines the two, except for the complicated MDP models. In this paper, we first combine stubborn and selfish mining attacks, i.e., construct a strategy where the attacker acts stubborn until its private branch reaches a certain length and then switches to act selfish. We provide the optimal stubbornness for each parameter regime. Next, we provide the maximum stubbornness that is still more profitable than honest mining and argue a connection between the level of stubbornness and the $k$-confirmation rule. We show that, at each attack cycle, if the level of stubbornness is higher than $k$, there is a risk of double-spending which comes at no-cost to the adversary. The result can be seen as a guide for picking $k$ in the $k$-confirmation rule in a blockchain design. At each cycle, for a given stubbornness level, we rigorously formulate how great the risk of double-spending is. We provide the minimum double-spend value needed for an attack to be profitable in the regimes where the scheme is less profitable than honest mining. We further modify the attack in the stubborn regime in order to conceal the attack and increase the double-spending probability. Finally, we evaluate the results and provide the optimal and the maximum stubbornness levels for each parameter regime as well as the revenue. As a case study, with Bitcoin’s $k=6$ block confirmation rule, we evaluate the revenue and double-spending risk of the attacks for each pool parameter.
@@ 1 Introduction
After Nakamoto’s seminal paper[1] that laid the foundations of blockchains aiming for large-scale consensus, there have been numerous works studying the related security guarantees and flaws [2]. The private attack, investigated initially in a non-rigorous manner in the Bitcoin whitepaper [1], aims to replace a transaction in the blockchain ledger with a conflicting transaction, allowing the attacker to double-spend its funds. Rosenfeld [3] showed rigorously that the private attack has exponentially decaying probability of success with time under the longest chain protocol. More specifically, as time passes, more public blocks are mined on top of a transaction that is being attacked, and it becomes less likely for a competing private chain to overtake the longest chain and undo the transaction in question.
@@
The celebrated selfish mining attack of Eyal and Sirer [4] exposed another vulnerability of the longest chain protocol by showing that it is not incentive-compatible. More specifically, [4] showed that a pool of miners, who constitute $\alpha$ proportion of the total hashrate in the network, can increase their revenue ratio $\rho$ by keeping the blocks they have mined private for a time according to the selfish mining algorithm, resulting in $\rho>\alpha$. This algorithm, analyzed via the use of a Markov chain, aims to waste the resources of non-pool miners by replacing their blocks with the private blocks of the pool and hence stealing their revenues. Moreover, the extra revenue of selfish mining, i.e., $\rho-\alpha$, grows even further as the pool size increases, which attracts the rational miners to join the pool, eventually overtaking the chain completely.
@@
Later on, block withholding attacks and their effects on the revenue have gained further attention and been studied extensively [5, 6, 7, 8]. Sapirshtein et. al. [6] use Markov Decision Processes (MDP) to provide $\epsilon$-optimal policies for selfish mining attacks that result in more profit than the algorithm of [4]. In a concurrent work, Nayak et. al. [7] provide an in-depth look at variations of selfish mining attacks, including the-now-famous stubborn mining, and analyze their revenues using Monte Carlo simulation techniques. Further, Gervais et. al. [8] create an extended MDP problem where the model takes into account additional parameters such as the confirmation rule, network propagation dynamics, mining costs and eclipse attack. The authors of [8] further modify the model to consider the profitability of the double-spending attack via the MDP and investigate the impacts of the parameters on the optimal policies. Similar to the study of [8], [9] builds an MDP problem to investigate the profitability of the double-spending attack by slightly modifying the MDP model of [6] that replaces the rewards of override blocks by double-spend blocks defined therein. The authors in [9] further modify the MDP to investigate the strategies that optimally combine selfish mining attacks with double-spend attacks assuming a constant relation between the profits from the two attacks. The authors of [10] investigate a novel efficient approach to solve average reward ratio MDP of [6] using probabilistic termination methods. Other generalizations of the MDP model also exist that analyze selfish mining in different proof systems and protocols [11, 12, 13, 14, 15]. Another line of research [16, 17] investigates the profitability of double-spending attacks using stochastic methods.
@@
An important feature of selfish mining attack is the network influence of the pool, which determines the proportion of non-pool miners who favor the adversarial chain in case of fork-ties. This parameter, which is called $\gamma$, provides a threshold for the profitability of selfish mining, i.e., the minimum pool size needed for the selfish mining attack to be profitable. In other words, $\gamma$ determines the minimum $\alpha$ needed such that $\rho(\alpha,\gamma)>\alpha$. The authors of [4] argue that the current bitcoin protocol, which picks the first chain seen by a miner in case of a tie, is vulnerable as the cost of increasing network influence in this case is minimal for the adversary. This in turn, resulted in search for new consensus policies that increase the threshold and the security [4, 18, 19, 20]. Other studies provide methods to effectively calculate network influence $\gamma$ under various network models [21, 22]. We refer the interested reader to the survey in [23] for a further read on selfish mining and double-spending attacks and related defense mechanisms.
@@ 1.1 Our Contributions
In this paper, for the first time in the literature, we provide a rigorous and explicit stochastic analysis of the combination of the two celebrated attacks: double-spending and selfish mining. To do so, we define a selfish mining attack strategy, that not only tracks the length difference between the pool’s private chain and the public chain but also its length. Tracking the length of the private chain allows us to determine whether the attack can result in double-spending or not in addition to the revenue steal from the honest miners. Under the $k$-confirmation rule, we argue that if the private branch of a selfish miner reaches length $k+1$, then the double-spending comes at no-cost. Thus, the question becomes whether it is profitable in terms of mining revenues to keep the private branch until it reaches length $k+1$. In other words, we investigate whether the new mining strategy is profitable, which is essentially a mix of selfish mining and stubborn mining that we call $L$-stubborn mining that acts stubborn until the private chain reaches length $L$, then acts selfish, i.e., releases the chain when the difference between the private and the public branches reduces to $1$.
@@
For each $\alpha$ and $\gamma$, we find the optimal $L^{*}$, i.e., the best strategy, under the condition that the adversary does not do any trailing behind, i.e., it accepts the honest branch whenever it falls behind. Similarly, we find largest $L$, that we call $\bar{L}$, such that $\rho>\alpha$, i.e., maximum stubbornness that still results in higher revenues than honest mining. Next, we argue that if $\bar{L}>k$, then double-spending comes at no-cost. Hence, the results not only give a new understanding of selfish and stubborn mining but also provide parameter regimes of $\alpha$ and $\gamma$ where transactions are at risk from the perspective of selfish mining. It is already known that when $\alpha\geq 0.5$ no transaction is safe. With the results we provide, for example for Bitcoin which follows $k$-confirmation rule with $k=6$, even if $\gamma=0$, every transaction is at risk of double-spend for $\alpha>0.409$. This, however, does not mean, that the transaction is going to be replaced with a conflicting one without fail, instead, it means, that trying the double-spending attack with no-trailing comes at no-cost to the adversary. Moreover, the attacks have three possible outcomes from the perspective of the transaction:
  1. 1.
    (Service) The original transaction is confirmed and stays in the longest chain in the long run.
  2. 2.
    (Move-Funds) The original transaction is mined into a block, however, it is replaced by a conflicting transaction before being confirmed.
  3. 3.
    (Double-Spend) The original transaction is confirmed and later replaced by a conflicting transaction.
We note that if the conflicting transaction is a transaction, where adversary moves its money between sybil accounts, move-funds event can be associated with a second chance that adversary gets at double-spending. Moreover, for a blockchain with finite block size, it decreases the effective throughput. For each $\alpha$ and $\gamma$, we provide the success probability of each event above, which essentially shows how great that risk is. For example, the double-spending risk (i.e., the double-spending probability) for $k=6$, $\alpha=0.41$ and $\gamma=0$ is $0.092$. For $k=6$ and $\gamma=0$, as $\alpha$ approaches $0.5$, the double-spending risk approaches $0.21$. On the other hand, for $k=6$, $\alpha=0.41$ and $\gamma=1$, the double spending risk is $0.002$, whereas service probability is $0.59$. Finally, we note that the scheme we provide assumes no pre-mining, hence, the security threats could be much worse under settings with pre-mining.
We also provide a modified version of $L$-stubborn mining attack, which we call $S$-stealth mining, that is more similar to the classical private attack, in the sense that it does not expose itself by matching blocks and increases the success probability of double-spending attack with no-trailing at the cost of revenue and move-funds event. We analyze the properties of $S$-stealth mining as we do for the $L$-stubborn mining attack.
@@
For both $L$-stubborn mining and $S$-stealth mining attacks, we provide a new metric for revenue calculation, where the extra revenue from the double-spending attack is taken into consideration in addition to the usual block rewards. Further, we also provide minimum transaction value for double-spending attacks that makes sure that the overall average revenue is larger than the honest mining protocol, when the usual revenue is less than the honest mining protocol.
In our analysis, we resort to the Bertrand’s ballot problem[24], Bailey’s number[25] and Catalan numbers[26], which have appeared explicitly in the blockchain literature in [27] where Chen et. al. calculate the probability that a chain wins a fork race given different mining probabilities depending on which fork is leading. Although Chen et. al. correctly argue that celebrated blockchain problems such as Nakamoto’s security problem [1] and selfish mining phenomenon [4] can be cast as modified ballot problems, they only provide the probabilities of each fork being ahead and their asymptotic values. Catalan numbers further appeared in [28, 29] where Grunspan and Marco make use of them to calculate the revenue of the stubborn mining [7] in closed form.
@@
Our analysis considers average reward ratio MDP of the [6, 10] for no-trail stubbornness and we provide a novel algorithm to solve the problem, which outputs the optimal values in finite steps. Our approach can be a basis to solve the general MDP problem, i.e., with trail stubbornness, directly, which we leave for future research.
We summarize the contributions of our paper as follows:
@@
  • Given pool size $\alpha$ and pool’s network influence $\gamma$, we find the optimal level of no-trail stubbornness $L$ and provide the associated revenue.
@@
Different from [8, 9], we provide a simple formula to calculate revenues, double-spend risks and optimal attacks. As we do not deal with MDPs, the numerical results in Section 6 are extremely easy-to-analyze and provide a great insight at first sight. Moreover, the strategy we provide is easy-to-follow compared to complicated prescriptions of MDPs.
@@ 2.1 Catalan Numbers and Dyck Words
Clearly $P[n,m]=0$ if $m>n$ and it is easy to verify that $P[n,0]=1$ and $P[n,1]=n$. Thus, (4) gives a straightforward formula for finding $P[n,m]$ in $\mathcal{O}(nm)$. Every Dyck word with length $2n$ has a suffix that starts with $1$ $X$ and $r$ $Y$’s, which implies
@@ 2.2 Bailey’s Number and Bertrand’s Ballot Problem
Note that, if a broad range of $P[n,m]$ is needed, it is better to start from $P[1,0]=1$ and fill the matrix $P$ via (4).
The elements of $P$ matrix appear in [25] where Bailey investigates the sequences that contain $n$ $1$’s and $m$ $-1$’s such that the running sum is always non-negative. Bailey proved that the number of such sequences is equal to $P[n,m]$ provided in (6). Further, the diagoanal and antidiagonal (including all the below diagonal and antidiagonal) elements of matrix $P$ overlap with the columns of $A^{(1,2)}$ defined in [31] as Catalan-like numbers, which is not surprising since the recursion equations are similar.
@@ 2.3 Wald’s Identity and Doob’s Optional Stopping Time
Consider a Pre-Dyck word with $L$ $X$’s and $m<L$ $Y$’s. Append a random sequence of $X$’s (w.p. $\alpha$) and $Y$’s (w.p. $1-\alpha$) until the first time the number of $Y$’s becomes 1 less than the number of $X$’s. As this is a stopping time, by Wald’s identity, the expected length of such a sequence is $\frac{L-m-1}{1-2\alpha}$. As $L-m-1$ $Y$’s are required to reduce the gap to $1$, the rest of the characters in the string are split evenly between $X$ and $Y$. Thus, the final sequence contains $L+\frac{(L-m-1)\alpha}{1-2\alpha}$ $X$’s in expectation.
@@ 3 System Model
The setting we consider in this paper for the blockchain model is the same as the model considered in the seminal work on selfish mining by Eyal and Sirer[4]. We assume that there are three groups of miners. The first group can be considered as a single-entity which constitutes $\alpha=1-\beta$ fraction of the mining power in the system and acts adversarially, i.e., does not follow the longest-chain protocol and can decide to mine privately without sharing the blocks it has mined. The second and third groups together comprise the honest miners who follow the longest-chain protocol, i.e., mine on top of the longest-chain they have seen and release the newly mined blocks immediately. The former of the two constitutes $\gamma\beta$ fraction of the mining power is influenced by the adversarial group and accepts the adversarial chain in case of a tie, whereas the latter constitutes $(1-\gamma)\beta$ fraction of the mining power and accepts the honest chain in case of a tie. The reason for such a partition and in which real-world scenarios this can arise is explained in [4].
@@
We only consider the blocks that contain a valid PoW (Proof-of-Work) and assume that the inter-arrival times between blocks follow an exponential distribution. Further, block propagation delays are assumed to be negligible, hence no blocks are mined while a new block is being propagated through the network. As a result, each mined block belongs to one of the three groups with a probability proportional to their mining power. Moreover, the blockchain protocol updates the puzzle difficulty threshold so that the average inter-arrival time between blocks is constant. Each block rewards its miner a predetermined amount of coins, i.e., coinbase, which we assume is constant for the foreseeable future. If each miner acts honestly, the coinbase reward they get is proportional to their mining power in the network in the limit. However, adversarial strategies which deviate from the longest-chain protocol result in rewards disproportional to the mining powers. We refer to the adversarial strategies that aim to increase their proportion of the coinbase rewards as selfish mining strategies. Although these selfish mining strategies initially decrease the rewards, they can be seen as an attack on the block interval adjustment [32], which in turn increases the rewards in the long run.
@@
A transaction $tx$ in this system is considered to be confirmed according to a $k$-deep confirmation rule if a block that contains $tx$ has $k-1$ blocks mined on top of it and is part of the longest chain available to the honest miners. Double-spending refers to the event that $tx$ is confirmed according to the $k$-deep confirmation rule and a conflicting transaction $tx^{\prime}$ replaces $tx$ in the longest chain and becomes $k$-deep.

### v2

@@ Abstract
Conventional double-spending attack models ignore the revenue losses stemming from the orphan blocks. On the other hand, selfish mining literature usually ignores the chance of the attacker to double-spend at no-cost in each attack cycle. In this paper, we give a rigorous stochastic analysis of an attack where the goal of the adversary is to double-spend while mining selfishly. To do so, we first combine stubborn and selfish mining attacks, i.e., construct a strategy where the attacker acts stubborn until its private branch reaches a certain length and then switches to act selfish. We provide the optimal stubbornness for each parameter regime. Next, we provide the maximum stubbornness that is still more profitable than honest mining and argue a connection between the level of stubbornness and the $k$-confirmation rule. We show that, at each attack cycle, if the level of stubbornness is higher than $k$, the adversary gets a free shot at double-spending. At each cycle, for a given stubbornness level, we rigorously formulate how great the probability of double-spending is. We further modify the attack in the stubborn regime in order to conceal the attack and increase the double-spending probability.
@@ 1 Introduction
After Nakamoto’s seminal paper[1] that laid the foundations of blockchains aiming for large-scale consensus, there have been numerous works studying the related security guarantees and flaws [2]. The private attack, investigated initially in the Bitcoin whitepaper [1], aims to replace a transaction in the blockchain ledger with a conflicting transaction, allowing the attacker to double-spend its funds. Rosenfeld [3] showed rigorously that the private attack has exponentially decaying probability of success with time under the longest chain protocol. More specifically, as time passes, more public blocks are mined on top of a transaction that is being attacked, and it becomes less likely for a competing private chain to overtake the longest chain and undo the transaction in question.
@@
The celebrated selfish mining attack of Eyal and Sirer [4] exposed another vulnerability of the longest chain protocol by showing that it is not incentive-compatible. More specifically, [4] showed that a pool of miners, who constitute $\alpha$ fraction of the total hashrate in the network, can achieve a revenue ratio $\rho$, that is more than their fair share $\alpha$, i.e., the fraction of the hashrate they posses in the system. This is achieved by keeping the blocks they have mined private for a while according to the selfish mining algorithm and releasing them later to replace the blocks mined by honest miners, resulting in $\rho>\alpha$ in the long run. Moreover, the extra revenue ratio of selfish mining, i.e., $\rho-\alpha$, grows even further as the pool size increases, which attracts the rational miners to join the pool, eventually overtaking the chain completely.
@@
Later on, block withholding attacks and their effects on the revenue ratio have gained further attention and been studied extensively [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Sapirshtein et. al. [5] use Markov Decision Processes (MDP) to provide $\epsilon$-optimal policies for selfish mining attacks that result in higher revenue ratio than the algorithm of [4]. In a concurrent work, Nayak et. al. [6] provide an in-depth look at variations of selfish mining attacks, including the-now-famous stubborn mining, and analyze their revenue ratio using Monte Carlo simulation techniques. Further, Gervais et. al. [7] create an extended MDP problem where the model takes into account additional parameters such as the confirmation rule, network propagation dynamics, mining costs and eclipse attack. The authors of [7] further modify the model to consider the profitability of the double-spending attack via the MDP and investigate the impacts of the parameters on the optimal policies. Similar to the study of [7], [8] builds an MDP problem to investigate the profitability of the double-spending attack by slightly modifying the MDP model of [5] that replaces the rewards of override blocks by double-spend blocks defined therein. The authors in [8] further modify the MDP to investigate the strategies that optimally combine selfish mining attacks with double-spend attacks assuming a constant relation between the profits from the two attacks. The authors of [9] investigate a novel efficient approach to solve average reward ratio MDP of [5] using probabilistic termination methods. Other generalizations of the MDP model also exist that analyze selfish mining in different proof systems and protocols [10, 11, 12, 13, 14, 15]. Selfish mining strategies can further be combined with block withholding attacks on pools to increase revenue ratio [20, 21, 22]. Another line of research [16, 17] investigates the profitability of double-spending attacks using stochastic methods. Some of the recent works also focus on the profitability over time and use alternative metrics such as revenue per time or chain progress [23, 24, 25, 26].
@@
An important feature of selfish mining attack is the network influence of the pool, which determines the fraction of non-pool miners who favor the adversarial chain in case of fork-ties. This parameter, which is called $\gamma$, provides a threshold for the profitability of selfish mining, i.e., the minimum pool size needed for the selfish mining attack to be profitable. In other words, $\gamma$ determines the minimum $\alpha$ needed such that $\rho(\alpha,\gamma)>\alpha$. The authors of [4] argue that the current Bitcoin protocol, which picks the first chain seen by a miner in case of a tie, is vulnerable as the cost of increasing network influence in this case is minimal for the adversary. This in turn, resulted in search for new consensus policies that increase the threshold and the security [4, 27, 28, 29, 10]. Other studies provide methods to effectively calculate network influence $\gamma$ under various network models [30, 31]. We refer the interested reader to the survey in [32] for a further read on selfish mining and double-spending attacks and related defense mechanisms.
@@ 1.1 Our Contributions
Double-spending attacks in blockchain literature usually assume an adversary that does not focus on coinbase rewards and purely tries to undermine the consistency of the protocol. Such an analysis that ignores the adversarial revenue loss is usually not realistic. Hence, in this paper, for the first time in the literature, we provide a rigorous and explicit stochastic analysis of the combination of the two celebrated attacks: double-spending and selfish mining. Following the attack prescription we provide, gives the adversary repeated chances at double-spending without incurring revenue losses. To do so, we define a selfish mining attack strategy, called $L$-stubborn mining, that not only tracks the length difference between the pool’s private chain and the public chain but also its length. Tracking the length of the private chain allows us to determine whether the attack can result in double-spending or not in addition to the revenue steal from the honest miners. Essentially, $L$-stubborn mining, parametrized by the length of the private chain $L$, is a generalization of existing mining strategies such as honest mining ($L=1$), selfish mining ($L=2$), equal fork stubborn mining ($L=\infty$) where $L$ represents the stubbornness level of the attacker pool. Thus, an optimization of $L$ results in optimal revenue ratios $\rho_{L^{*}}$ among the no-trail stubborn strategies [6].
@@
More specifically, for each $\alpha$ and $\gamma$, we find the optimal $L^{*}$, i.e., the best strategy, under the condition that the adversary does not do any trailing behind, i.e., it accepts the honest branch whenever it falls behind. Similarly, we find largest $L$, that we call $\bar{L}$, such that $\rho>\alpha$, i.e., maximum stubbornness that still results in higher revenue ratio than honest mining strategy. Next, under the $k$-confirmation rule, we argue that if $\bar{L}>k$, then double-spending comes at no-cost. Hence, the results not only give a new understanding of selfish and stubborn mining but also provide parameter regimes of $\alpha$ and $\gamma$ where transactions are at risk from the perspective of double-spending. It is already known that when $\alpha\geq 0.5$ no transaction is safe. With the results we provide, for example for Bitcoin which follows $k$-confirmation rule with $k=6$, even if $\gamma=0$, every transaction is at risk of double-spend for $\alpha>0.409$. This, however, does not mean, that the transaction is going to be replaced with a conflicting one without fail, instead, it means, that trying the double-spending attack with no-trailing comes at no-cost to the adversary.
@@
We also provide a modified version of $L$-stubborn mining attack, which we call $S$-stealth mining, that is more similar to the classical private attack, in the sense that it does not expose itself by matching blocks and increases the success probability of double-spending attack at the cost of revenue. We analyze the properties of $S$-stealth mining as we do for the $L$-stubborn mining attack. For both $L$-stubborn mining and $S$-stealth mining attacks, we provide a new metric for revenue ratio calculation similar to the subversion gain defined in [10], where the extra revenue from the double-spending attack is taken into consideration in addition to the usual block rewards. Further, we also provide minimum transaction value for double-spending attacks that makes sure that the overall revenue ratio is larger than the honest mining strategy, when the usual revenue ratio is less than the honest mining strategy, similar to the subversion bounty defined in [10].
@@
Our analysis considers average reward ratio MDP of the [5, 9] for no-trail stubbornness and we provide a novel algorithm to solve the problem, which outputs the optimal values in finite steps. Our approach can be a basis to solve the general MDP problem, i.e., with trail stubbornness, directly, which we leave for future research. We summarize the contributions of our paper as follows:
@@
  • Given pool size $\alpha$ and pool’s network influence $\gamma$, we find the optimal level of no-trail stubbornness $L$ and provide the associated revenue ratio.
@@
Different from [7, 8], we provide a simple formula to calculate revenue ratios, double-spend risks and optimal attacks. As we do not deal with MDPs, the numerical results in Section 6 are extremely easy-to-analyze and provide a great insight at first sight. Moreover, the strategy we provide is easy-to-follow compared to complicated prescriptions of MDPs. Our results can directly be used to rigorously analyze optimal mining revenue per time, a recent focus of research topic [23, 24, 25, 26].
In our analysis, we resort to the Bertrand’s ballot problem[33], Bailey’s number[34] and Catalan numbers[35], which have appeared explicitly in the blockchain literature in [36] where Chen et. al. calculate the probability that a chain wins a fork race given different mining probabilities depending on which fork is leading. Although Chen et. al. correctly argue that celebrated blockchain problems such as Nakamoto’s security problem [1] and selfish mining phenomenon [4] can be cast as modified ballot problems, they only provide the probabilities of each fork being ahead and their asymptotic values. Catalan numbers further appeared in [37, 38] where Grunspan and Marco make use of them to calculate the revenue ratio of the stubborn mining [6] in closed form and in [39] where Goffard refines the mathematical model of double-spending.
Our paper is structured as follows. In Section 2, we provide the essential probabilistic tools in a concise manner that are needed to understand the analysis provided in this paper. In Section 3, we define the system model, provide the attackers’ mining strategies and relevant definitions. In Section 4, we rigorously analyze the revenue ratio of the $L$-stubborn attack as well as the double-spending that comes with it. We do the same for $S$-stealth mining strategy in Section 5. In Section 6, we provide the relevant numerical results. Finally, in Section 7, we discuss the severity and potential implications of the attack as well as possible extensions.
@@ 2.1 Catalan Numbers and Dyck Words
Clearly $P[n,m]=0$ if $m>n$ and it is easy to verify that $P[n,0]=1$ and $P[n,1]=n$. Thus, (4) gives a straightforward formula for finding $P[n,m]$ in $\mathcal{O}(nm)$. Every Dyck word with length $2n$ has a suffix that starts with 11 $X$ and $r$ $Y$’s, which implies
@@ 2.2 Bailey’s Number and Bertrand’s Ballot Problem
Note that, if a broad range of $P[n,m]$ is needed, it is better to start from $P[1,0]=1$ and fill the matrix $P$ via (4). The elements of the matrix $P$ constitute Catalan’s triangle.
The elements of $P$ matrix appear in [34] where Bailey investigates the sequences that contain $n$ 11’s and $m$ 1-1’s such that the running sum is always non-negative. Bailey proved that the number of such sequences is equal to $P[n,m]$ provided in (6). Further, the diagoanal and antidiagonal (including all the below diagonal and antidiagonal) elements of matrix $P$ overlap with the columns of $A^{(1,2)}$ defined in [41] as Catalan-like numbers, which is not surprising since the recursion equations are similar.
@@ 2.3 Wald’s Identity and Doob’s Optional Stopping Time
Consider a Pre-Dyck word with $L$ $X$’s and $m<L$ $Y$’s. Append a random sequence of $X$’s (with probability $\alpha$) and $Y$’s (with probability $1-\alpha$) until the first time the number of $Y$’s becomes 1 less than the number of $X$’s. As this is a stopping time, by Wald’s identity, the expected length of such a sequence is $\frac{L-m-1}{1-2\alpha}$. As Lm1L-m-1 $Y$’s are required to reduce the gap to $1$, the rest of the characters in the string are split evenly between $X$ and $Y$. Thus, the final sequence contains $L+\frac{(L-m-1)\alpha}{1-2\alpha}$ $X$’s in expectation.
@@ 3 System Model
We assume a blockchain running under the Nakamoto consensus, i.e., PoW with longest chain protocol. The mining model we consider is the same as the model used in the seminal work on selfish mining by Eyal and Sirer[4] and the subsequent works on the topic [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. More specifically, we assume that there are three groups of miners. The first group can be considered as a single-entity which constitutes $\alpha=1-\beta$ fraction of the mining power in the system and acts adversarially, i.e., does not follow the longest-chain protocol and can decide to mine privately without sharing the blocks it has mined. In this paper, we call this entity as the adversary. The second and third groups together comprise the honest miners who follow the longest-chain protocol, i.e., mine on top of the longest-chain they have seen and release the newly mined blocks immediately. The former of the two constitutes $\gamma\beta$ fraction of the mining power is influenced by the adversarial group and accepts the adversarial chain in case of a tie, whereas the latter constitutes $(1-\gamma)\beta$ fraction of the mining power and accepts the honest chain in case of a tie. The reason for such a partition and in which real-world scenarios this can arise is explained in [4].
We only consider the blocks that contain a valid PoW (Proof-of-Work) and assume that the inter-arrival times between blocks follow an exponential distribution. When a block is mined by an honest miner, it is subject to a propagation delay. During this propagation delay, if the adversary releases a competing block at the same height, called matching, $\gamma$ fraction of the honest miners receive the adversarial block first and mine on top of it. On the other hand, we ignore the newly mined blocks during the propagation delay. As a result, each mined block belongs to one of the three groups with a probability proportional to their mining power.
@@
Moreover, the blockchain protocol updates the puzzle difficulty threshold so that the average inter-arrival time between blocks is constant. Each block rewards its miner a predetermined amount of coins, i.e., coinbase, which we assume is constant for the foreseeable future. If each miner acts honestly, the coinbase reward they get is proportional to their mining power in the network in the limit. However, adversarial strategies which deviate from the longest-chain protocol result in rewards disproportional to the mining powers. We refer to the adversarial strategies that aim to increase their fraction of the coinbase rewards as selfish mining strategies. Although these selfish mining strategies initially decrease the rewards, they can be seen as an attack on the block interval adjustment [42], which in turn increases the rewards in the long run.
@@
A transaction $tx$ in this system is considered to be confirmed according to a $k$-deep confirmation rule if a block that contains $tx$ has k1k-1 blocks mined on top of it and is part of the longest chain available to the honest miners. Double-spending refers to the event that $tx$ is confirmed according to the $k$-deep confirmation rule and a conflicting transaction $tx^{\prime}$ replaces $tx$ in the longest chain and becomes $k$-deep.
3.1 Propagation Delays
We note that, from the perspective of the propagation delays, the model described here has been considered in the majority of the selfish mining literature and is still the convention. Essentially, if blocks can be mined during propagation delays, the analysis has to keep track of multiple (potentially infinite) different honest miner views that evolve as time goes by. Hence, in the selfish mining literature, the assumption of no block mining during propagation delays has been adopted for tractability purposes. On the other hand, studies on double-spending analysis, such as [43, 44, 45], consider the propagation delays, because from the perspective of the consistency guarantees, worst-case scenarios are relevant. Hence, these studies usually assume that the adversary controls the propagation delays as long as each propagation delay is lower than a maximum $\Delta$ and ties are broken in its favor. As a result, these analyses usually focus on a single worst-case honest miner view [44, Section III.A, weakest possible view]. However, such a situation is extremely pessimistic as our goal is not to give worst-case analysis of revenue ratios in selfish mining. We also note that, optimal selfish mining analysis has been done previously in such a pessimistic setting [46, Lemma 4].
Further, under Nakamoto consensus, given the block creation rate $\lambda$, if $\lambda\Delta$ is not small, the consistency of the system at threat and selfish mining is a less relevant concern. In such systems, the research focuses on different consensus models such as [47, 48, 49]. On the other hand, as long as $\lambda\Delta$ is small, it can be shown that the probability of more than $2$ blocks being mined within a $\Delta$ interval is negligible [50, 31, 51]. Thus, in the realistic scenario of one block being mined during the propagation delay of a block, during forks, the honest network divides between two branches. As a result, in the long run, the parameter $\gamma$ used in selfish mining literature is still a relevant and reasonable assumption. Further, in [50, Section V] the authors give a formula to estimate $\gamma$ given propagation delays between mining entities and validate the estimation with network simulations.
Essentially, as things become immediately untractable when one considers multiple miners’ views during network delays, the system model with $\alpha$ and $\gamma$ serve as tractable and efficient parameters to model selfish mining attacks in a simplified manner that still roughly translates to a system with propagation delays where $\gamma$ is the main parameter of interest from the perspective of the delays and fork conflicts and should be estimated based on the observed delays. Here, note that, propagation delays also affect the apparent hashrate of a mining entity. In other words, the apparent hashrate might deviate from the fraction of computational power the entity possesses due to the propagation delays and network structure even if everyone follows honest mining strategy. Notice that, propagation delays result in forks and the block published first does not always win the fork race. Hence, one can also think of $\alpha$ as the apparent hash rate instead of the fraction of computational power. For example, studies such as [51] consider different propagation delays between different miners to evaluate fairness, which is roughly defined as the difference between the fraction of computational power and the apparent hashrate a mining entity has. We believe that modeling such a system to estimate $\alpha$ and $\gamma$ values used in this paper are possible however out of scope of this paper. For further details about the propagation delays and their impact on fork races and fairness, we refer the reader to [50, 31, 51].